Lac Operon IPOZY v3 In the genotype presented (genomic plus plasmid genes), which of the following statements is consistent with the expression phenotype? genome: I+P+ O- Z+ Y- plasmid: I+ P- O+ Z- Y+

  • constitutive expression of β-gal and permease induced in the presence of lactose

  • constitutive expression of β-gal only

  • constitutive expression of β-gal and permease

  • β-gal and permease induced in the presence of lactose

  • constitutive expression of permease and β-gal induced in the presence of lactose
Answer #1

1. Answer: Option 5 is correct
Explanation:
The Lac operon is a negatively regulated inducible operon.
It is expressed only in the presence of lactose provided that glucose is absent.
Genotype = I+P+O-Z+Y-/I+P-O+Z-Y+
Genomic copy = I+P+O-Z+Y-
Since it is O-, Z+ (Beta-gal) is expressed in the presence of lactose.

Plasmid copy = I+P-O+Z-Y+
Since it is P-, Y+ (Permease) is never expressed.

= Inducible expression of LacZ and constitutive repression of LacY


A positive rating is highly appreciated. Thank you